OpenStudy (anonymous):

What is the approximate area of the shaded region? A. 104.86 cm² B. 87.47 cm² C. 27.02 cm² D. 10.54 cm²

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

Area of the square = 7² = 49 Area of the circle = (3.14)(3.5)² ≈ 38.47 Shaded regions: Area of Square - area of circle = 49 - 38.47 = 10.53

OpenStudy (anonymous):

3.5= radius, b/c 7 is diameter and 3.5 is a half of that. 3.14 is an approximate value f pi. it's D:) HAVE A GOOD DAY !

OpenStudy (anonymous):

thnx @student_basil

OpenStudy (anonymous):

You welcome !